# Controlled Low-Rank Adaptation with Subspace Regularization for Continued Training on Large Language Models

### v1

@@ 5.3 Understanding Capacity-Forgetting Balancing(RQ3)
Measuring Model Capacity and Degree of Forgetting
Consider that catastrophic forgetting primarily arises from output changes caused by parameter updating, the greater the impact of these updates, the more severe the catastrophic forgetting may be. We measure the degree of forgetting with the relative scale of output change in the parameter level, to be specific, for updating matrix $\Delta W$, with input $x$, the relative scale of output change (denoted as $\mathrm{F}$) is defined as
$\mathrm{F}(\Delta W,x)=\frac{||\Delta Wx||}{||x||}$ (3)
We highlight the role of $x$ in the measurement of “F”, as it reflects the real world use case. Specifically, we sample 100 text data from test set and collect input $x$ for each parameter from the model forward pass.
Results and Analysis
The “reference” row is computed using the LoRA trained model, noting the output scale of original parameter $W$ instead of $\Delta W$. Compared with “reference” and “LoRA”, the difference of “F” is not very far, suggesting that LoRA training indeed introduces significant output change, thus still prone to catastrophic forgetting.
For “MiLoRA”, although intuitively promising, without effective control during training, it did not mitigate catastrophic forgetting, as evidenced by both downstream evaluations (in section 5.1) and the similar “F” and $\Delta W$ with LoRA.
For LoRA with lower rank (r8 and r16), after training, with $||\Delta W||$ indicates the reduction of capacity, the scale of output change does not show a decrease. Although theoretically, reducing the rank of the update matrix can increase the dimension of the null space and help to reduce the scale of output change, results not show this case. This suggests that altering $r$ may not a effctive way to alter forgetting.
For LoRA-L2, “F” indicates that it indeed mitigate forgetting, but in a large cost of capacity, demonstrated by the very small $||\Delta W||$.
For our proposed CLoRA, “F” shows a significantly reduce the scale of output change, while a relatively larger $||\Delta W||$ is maintained. This indicates that CLoRA minimizes catastrophic forgetting caused by large updates while having a subtle impact on model capacity. Thus we answer RQ3 that CLoRA performs effectively on capacity-forgetting balancing.
Method $||\Delta W||$ $\mathrm{F}$
@@
Table 5: Measuing of model updating capacity($||\Delta W||$, larger for more capacity) and degree of forgetting($\mathrm{F}$, lower for less forgetting) for trained models.
@@ 6 Conclusion
In this paper, we introduce Controlled Low-Rank Adaptation(CLoRA), a simple yet effective parameter-efficient finetuning method for LLMs that mitigates catastrophic forgetting. We investigate the effectiveness of CLoRA on commonsense reasoning and math reasoning tasks. Experiment results demonstrate that CLoRA consistently outperforms previous LoRA-based methods on in-domain evaluations, highlighting the effectiveness of CLoRA as a parameter-efficient finetuning method. Additionally, out-domain evaluations demonstrate that CLoRA exhibits strong capabilities in mitigating catastrophic forgetting. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
@@ 7 Limitations
There are still several limitations that we reserve for future work: 1) While we have noted that our proposed CLoRA would benefit from the dedicated selection of regularization matrix to retain specific model abilities, we did not go further into this aspect. In this paper, we only explored the SVD initialization variant alongside the basic random initialization. 2) We delegate the measurement of model capacity and degree of forgetting to simple measurement of scale. Although these measurements reveal significant differences between CLoRA and previous works, we believe that further investigation would aid in the design of methods with stronger capacity-forgetting balancing capability.
Appendix A Detailed Experiment Setups
@@ A.2 Computation Environment
Appendix B Full Experiment Results
B.1 Full Results on Commonsense Finetuning
@@
We report the full results that we conducted on in-domain evaluation(Table 7) and out-domain evaluation(Table 8) for commonsense reasoning finetuning. All models use LoRA rank $r$ of 32 unless specified otherwise.

### v2

@@ 5.1.6 Understanding Capacity-Forgetting Balancing(RQ3)
Method $||\Delta W||$ $\mathbb{F}$
@@
Table 4: Measuring model updating capacity($||\Delta W||$, larger for more capacity) and degree of forgetting($\mathbb{F}$, lower for less forgetting) for trained models.
Measuring Model Capacity and Degree of Forgetting
Consider that catastrophic forgetting primarily arises from output changes caused by parameter updating, the greater the impact of these updates, the more severe the catastrophic forgetting may be. We measure the degree of forgetting with the relative scale of output change in the parameter level, to be specific, for updating matrix $\Delta W$, with input $x$, the relative scale of output change (denoted as $\mathbb{F}$) is defined as
$\mathbb{F}(\Delta W,x)=\frac{||\Delta Wx||}{||x||}$ (3)
We highlight the role of $x$ in the measurement of $\mathbb{F}$, as it reflects the real world case. Specifically, we sample 100 text data from test set and collect input $x$ for each parameter from the model forward pass.
Results and Analysis
The “reference” row is computed using the LoRA trained model, noting the output scale of original parameter $W$ instead of $\Delta W$. Compared with “reference” and LoRA, the difference of $\mathbb{F}$ is not far, suggesting that LoRA training indeed introduces significant output change, thus still prone to catastrophic forgetting.
For MiLoRA, although intuitively promising, without effective control during training, it did not mitigate catastrophic forgetting, as evidenced by both downstream evaluations (Table 2) and the similar $\mathbb{F}$ and $\Delta W$ with LoRA.
For LoRA with lower rank (r8/16), after training, with $||\Delta W||$ indicates the reduction of capacity, $\mathbb{F}$ does not show a decrease. Although theoretically, reducing the rank of the update matrix can increase the dimension of the null space and help to reduce the scale of output change, results not show this case. This suggests that altering $r$ may not a effctive way to alter forgetting.
For LoRA-L2, $\mathbb{F}$ indicates that it indeed mitigate forgetting, but in a large cost of capacity, demonstrated by the very small $||\Delta W||$.
For our proposed CLoRA, $\mathbb{F}$ shows a significantly reduce the scale of output change, while a relatively larger $||\Delta W||$ is maintained. This indicates that CLoRA minimizes catastrophic forgetting caused by large updates while having a subtle impact on model capacity. Thus we answer RQ3 that CLoRA performs effectively on capacity-forgetting balancing.
Standard CL Benchmark Large Number of Tasks
Method Order-1 Order-2 Order-3 avg. Order-4 Order-5 Order-6 avg.
SeqFT 18.9 24.9 41.7 28.5 7.4 7.4 7.5 7.4
SeqLoRA 44.6 32.7 53.7 43.7 2.3 0.6 1.9 1.6
IncLoRA 66 64.9 68.3 66.4 63.3 58.5 61.7 61.2
Replay 55.2 56.9 61.3 57.8 55 54.6 53.1 54.2
EWC 48.7 47.7 54.5 50.3 45.3 44.5 45.6 45.1
LwF 54.4 53.1 49.6 52.3 50.1 43.1 47.4 46.9
L2P 60.3 61.7 61.1 60.7 57.5 53.8 56.9 56.1
LFPT5 67.6 72.6 77.9 72.7 70.4 68.2 69.1 69.2
O-LoRA 75.4 75.7 76.3 75.8 72.3 64.8 71.6 69.6
CLoRA 79.7 79.1 78.2 79.0 70.7 65.6 68.2 68.1
PerTaskFT 70.0 70.0 70.0 70.0 78.1 78.1 78.1 78.1
MTL 80.0 80.0 80.0 80.0 76.5 76.5 76.5 76.5
Table 5: Results on two CL benchmarks with T5-large base model. Averaged accuracy after training on the last task is reported. Bold font indicates the highest performance across all compared CL methods.
5.2 Continual Learning
5.2.1 Experimental Setup
To demonstrate the effectiveness of CLoRA for continual learning(CL) setting, we conduct experiments on standard CL benchmark and more challenging large number of tasks benchmark, following the experiment setup of O-LoRAWang et al. (2023b).
Datasets and Tasks
The standard CL benchmark consists of five text classification datasetsZhang et al. (2015). The large number of tasks benchmark consists of 15 datasets Razdaibiedina et al. (2023), include tasks for natural language understanding and text classification. Task samples follows previous work Wang et al. (2023b). Details for tasks are listed in Appendix B.
Comparison Methods
We compare CLoRA with normal finetuning baselines and previous CL methods. We include non CL results that train separate model for each task (PerTaskFT) and multi-task learning (MTL) as reference.
  • Normal Finetuing baselines include sequentially training on same parameter space with full parameter finetune (SeqFT) and LoRA (SeqLoRA), and incremental learning of new LoRA parameters on a sequential series of tasks.
  • Continual Learning methods include data-based methods Replay; architecture-based methods L2PWang et al. (2022), LFPT5Qin and Joty (2022)O-LoRAWang et al. (2023b); and learning-based methods EWCKirkpatrick et al. (2017), LwFLeibe et al. (2016). Details for these methods are listed in Appendix B.
Experimental Configuration
Following O-LoRAWang et al. (2023b), we use T5-large as base model, and finetune on each task with specified order(Appendix B). We train each task with one epoch, with constant learning rate of 1e-3, batch size of 64, dropout rate of 0.1, weight decay rate of 0, and LoRA dim $r$ of 8. CLoRA regularization matrix size $k$ is set to 256.
5.2.2 Results and Analysis
We report the results in Table 5. Results demonstrate that CLoRA outperforms all comparision methods, include the most related strong baseline O-LoRA, with a notable margin in the stadard CL benchmark. We attribute this to the advantage of CLoRA toward O-LoRA: 1. CLoRA helps learning in the first finetuning stage while O-LoRA not; 2. CLoRA can independently alter $k$ for balancing learning and forgetting, while “$k$” equivalent in O-LoRA is restrained by LoRA $r$.
In the large number of tasks benchmark, CLoRA performs near the strong baseline O-LoRA and LFPT5. We note that vanilla CLoRA with random regularization matrix is a learning-based method, without machanism for isolating finetuning tasks.
@@ 6 Conclusion
In this paper, we introduce Controlled Low-Rank Adaptation(CLoRA), a simple yet effective parameter-efficient finetuning method for LLMs that mitigates catastrophic forgetting. We investigate the effectiveness of CLoRA on both one-stage LLM finetuning and continual learning settings. Experiment results demonstrate the effectiveness of CLoRA as a parameter-efficient finetuning method with catastrophic forgetting mitigating. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
@@ 7 Limitations
There are still several limitations that we reserve for future work: 1) We use the simplest random initialization for regularization matrix, insight for more dedicated choice may benifit CLoRA learning, such as combine CLoRA with architecture-based continual learning method(O-LoRA). 2) We delegate the measurement of model capacity and degree of forgetting to simple measurement of scale. Although these measurements reveal significant differences between CLoRA and previous works, we believe that further investigation would aid in the design of methods with stronger capacity-forgetting balancing capability.
Appendix A Experiment Details for One-Stage Finetuning
@@ A.2 Computation Environment
A.3 Additional CLoRA variants
We use the simplest random initialization for CLoRA regularization matrix in the main paper. Considering the idea of PiSSA and MiLoRA that explore the roles of singular value decomposition (SVD) components in LLM parameters, we adopt this intuition to initialize CLoRA regularization matrices from SVD. For a SVD decomposition of parameter $W=USV^{T}$ with rank $r$, where $W\in\mathbb{R}^{m\times n}$, $U\in\mathbb{R}^{m\times r}$, $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix, $V\in\mathbb{R}^{n\times r}$. For CLoRA updating $\Delta W=AB^{T}$, we initialize the regularization matrix $P_{A}\in\mathbb{R}^{m\times k}$ as $U[:,\mathbf{s}]$, $P_{B}\in\mathbb{R}^{n\times k}$ as $V[:,\mathbf{s}]$, where $\mathbf{s}$ is a list of selecting index with length $k$. We add two CLoRA variants as follows, and conduct experiments on commonsense reasoning setting,
A.4 Full Results on Commonsense Finetuning
@@
We report the full results that we conducted on in-domain evaluation(Table 7) and out-domain evaluation(Table 8) for commonsense reasoning finetuning. Results for LLaMA-3-8b are also included for CLoRA-random. All models use LoRA rank $r$ of 32 unless specified otherwise.
A.4.1 Analysis for different CLoRA variants
Results indicate that the choice of regularization matrix does influence the effectiveness of CLoRA, albeit not significantly. Generally, we recommend using random initialization (CLoRA-random) or initialization from major SVD components (CLoRA-major).
Appendix B Detailed Experiment Setups for Continual Learning
Dataset name Category Task Domain
Yelp CL Benchmark sentiment analysis Yelp reviews
Amazon CL Benchmark sentiment analysis Amazon reviews
DBpedia CL Benchmark topic classification Wikipedia
Yahoo CL Benchmark topic classification Yahoo Q&A
AG News CL Benchmark topic classification news
MNLI GLUE NLI various
QQP GLUE paragraph detection Quora
RTE GLUE NLI news, Wikipedia
SST-2 GLUE sentiment analysis movie reviews
WiC SuperGLUE word sense disambiguation lexical databases
CB SuperGLUE NLI various
COPA SuperGLUE QA blogs, encyclopedia
BoolQA SuperGLUE boolean QA Wikipedia
MultiRC SuperGLUE QA various
IMDB SuperGLUE sentiment analysis movie reviews
Table 9: Summary of datasets used in the continual learning setting.
Order Task Sequence
1 dbpedia $\rightarrow$ amazon $\rightarrow$ yahoo $\rightarrow$ ag
2 dbpedia $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo
3 yahoo $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ dbpedia
4 mnli $\rightarrow$ cb $\rightarrow$ wic $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ boolqa $\rightarrow$ rte $\rightarrow$ imdb
$\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ multirc $\rightarrow$ yahoo
5 multirc $\rightarrow$ boolqa $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte
$\rightarrow$ imdb $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ yahoo
6 yelp $\rightarrow$ amazon $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb
$\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ multirc $\rightarrow$ boolqa $\rightarrow$ wic
Table 10: Order of finetuning in the continual learning setting.
B.1 Dataset Details
We list the details of the datasets used in Table 9. Order of finetuning are listed in Table 10.
B.2 Computation Environment
All of our experiments are conducted on 1 NVIDIA GeForce RTX 3090 GPU. All methods for LoRA subsequents use Huggingface peft library, training is conducted using trainer in Huggingface transformers library, with DeepSpeed ZeRO intergration.
B.3 Comparision Methods
Here we provide details for continual learning baselines for our continual learning experiment setting.
  • Replay: data-based method that replay samples from old tasks when learning new tasks to avoid forgetting.
  • L2P: architecture-based method that uses the input to dynamically select and update prompts from the prompt pool in an instance-wise fashion.
  • LFPT5: architecture-based method that continuously train a soft prompt that simultaneously learns to solve the tasks and generate training samples for replay.
  • EWC: learning-based method that finetune the whole model with a regularization loss that prevents updating parameters that could interfere with previously learned tasks.
  • LwF: learning-based method that constrains the shared representation layer to be similar to its original state before learning the new task.
  • O-LoRA: architecture and learning-based method that prevent subsequent LoRA update interfere previous.